# AgentExpt: Automating AI Experiment Design with LLM-based Resource Retrieval Agent

### v1

@@
Abstract.
Large language model agents are becoming increasingly capable at web-centric tasks such as information retrieval, complex reasoning and online navigation. These emerging capabilities have given rise to surge research interests in developing LLM agents for facilitating scientific quest, which is often known as AI Scientist systems. One key application in AI research is to automate AI experiment design through agentic dataset and baseline retrieval, which have the potential to largely accelerate AI experiments and assist peer review. However, prior efforts suffer from limited data coverage, as recommendation datasets primarily harvest candidates from public portals and omit many datasets actually used in published papers, and from an overreliance on content similarity that biases models toward superficial similarity and overlooks experimental suitability. Harnessing collective perception embedded in the baseline and dataset citation network, we present a comprehensive framework for baseline and dataset recommendation. First, we design an automated data-collection pipeline that links roughly one hundred thousand accepted papers to the baselines and datasets they actually used. Second, we propose a collective perception enhanced retriever. To represent the position of each dataset or baseline within the scholarly network, it concatenates self-descriptions with aggregated citation contexts. To achieve efficient candidate recall, we finetune an embedding model on these representations. Finally, we develop a reasoning-augmented reranker that exacts interaction chains to construct explicit reasoning chains and finetunes a large language model to produce interpretable justifications and refined rankings. The dataset we curated covers 85% of the datasets and baselines used at top AI conferences over the past five years. On our dataset, the proposed method outperforms the strongest prior baseline with average gains of +5.85% in Recall@20, +8.30% in HitRate@5, and +7.90% in HitRate@10. Taken together, our results advance reliable, interpretable automation of experimental design.
baseline recommendation, dataset recommendation, citation contexts, reranking, LLM agents
Figure 1. Overview of the Research Problem:
1. Introduction
Recent advances in large language models (LLMs) have rapidly expanded their capabilities(Brown et al., 2020; Radford et al., 2019, 2018; Shang et al., 2024b, a; Li et al., 2025). In particular, LLM-powered agents have been widely applied to web-centric tasks, including web navigation and browsing (Zhou et al., 2023; Deng et al., 2023; Chezelles et al., 2024; Lù et al., 2024), personalized recommendation and user modeling (Zhang et al., 2024; Wang et al., 2023; Hao et al., 2025), and goal-oriented online shopping (Yao et al., 2022). At the same time, there is growing interest in leveraging LLM agents to gather information from the web for scientific tasks such as literature review and idea generation (Huang et al., 2025; Team et al., 2025). Building on this trend, end-to-end research frameworks such as Carl (Institute, 2025), Zochi (Intology, 2025), and the AI Scientist (Lu et al., 2024; Yamada et al., 2025) have been proposed to automate the research pipeline. However, these systems typically operate within narrow, pre-defined problem settings and do not require careful selection of baselines or datasets, which limits their generality and rigor. Meanwhile, the number of available models and datasets in machine learning has grown dramatically, making it increasingly important and challenging to recommend appropriate baselines and datasets for a new research idea, see fig 1.
Figure 2. Constructing chain-derived candidates and analysis. Left: From the interaction graph of papers, baselines and datasets, we extract interaction chains and aggregate the terminal items to form a chain-derived dataset/baseline pool. Right: (i) Recall (%) between a target paper’s actual baselines/datasets and candidates from each setting, and (ii) Precision (%) of overlapped items within the corresponding candidate pool. We evaluate three settings on both the baseline side and the dataset side: chain-derived top–100, same conference-derived top–100, embedding top- 100. Chain-derived candidates recover on average 60.14% of baselines and 78.61% of datasets while occupying 2.52% and 5.63% of the respective chain-derived pools, indicating that interaction chains provide a compact yet highly informative prior for selecting baselines and datasets.:
Research on dataset and baseline recommendation spans two components: datasets and methods. On the dataset side, RecBaselines2023 targets baseline recommendation within recommender systems (Ivanova et al., 2023), while DataFinder and DataHunter address dataset recommendation from problem statements or open retrieval (Viswanathan et al., 2023a; Färber and Leisinger, 2021b, c). On the method side, representative approaches include supervised text classification (Färber and Leisinger, 2021c, b), dense bi encoder retrieval (Viswanathan et al., 2023a), graph based modeling over papers and datasets (Altaf et al., 2019; Qayyum et al., 2025), and collaborative filtering over baseline co occurrences (Ivanova et al., 2023). Collectively, these strands establish initial tasks and evaluation setups. Despite progress, important gaps remain. On the data side, existing baseline recommendation dataset is limited to specific domains (Ivanova et al., 2023) and scales, while dataset recommendation benchmarks primarily harvest candidates from portals such as Papers with Code (Viswanathan et al., 2023a) or Wikidata (Färber and Leisinger, 2021c, b),omitting many datasets that are actually used in published papers. This limits coverage and reduces practical utility for real research workflows. On the method side, most systems (Färber and Leisinger, 2021b; Beltagy et al., 2019; Qayyum et al., 2025; Goyal et al., 2024; Gu et al., 2022) rely solely on first-person self-descriptions which biases models toward surface-level semantic similarity and yields items that appear topically related yet are ill-suited to the intended experimental setting. Moreover, prior work typically treats baseline and dataset recommendation independently, overlooking their synergy.
@@ 1. Introduction
Motivated by our day-to-day research practice, we typically begin from a set of closely related papers and follow their trails to the baselines and datasets they actually used, downloading artifacts directly rather than browsing portal listings. In choosing these components, we consider baselines and datasets jointly, since pairings that have been used together—or are methodologically compatible—are more likely to fit our intended evaluation setting. Moreover, when writing a paper we must justify these choices, situating a baseline or dataset relative to prior work. So the citation contexts constitute third-person signals that complement an asset’s first-person self-description. These observations motivate a unified formulation that links papers, baselines, and datasets and explicitly leverages community-level perception.
Building on this motivation, to build a massive baseline and dataset recommendation dataset that explicitly links research papers to the baselines and datasets they actually use, we design an automated data-collection pipeline that gathers roughly one hundred thousand accepted papers from the past decade across ten flagship AI venues. We convert PDFs to structured text, identify and parse experiment-centric sections, and apply LLM-assisted extraction to recognize true mentions of baseline methods and evaluation datasets. We then normalize naming variants, merge duplicates, and associate each entity with its introducing paper, yielding a high-precision mapping from papers to the experimental components they employ. On this dataset, to capture how the community actually positions and uses a baseline or dataset, we extract citation contexts as third-person perception signals that complement first-person self-descriptions. For each target (baseline or dataset), we summarize its citation contexts and represent it by concatenating its self-description with the aggregate community perception. To efficiently recall relevant candidates, we finetune an embedding model on these representations. To exploit baseline–dataset synergy(see fig 2) and refine the shortlist, we develop a reasoning-augmented reranking framework that extracts paper–baseline–paper-dataset and paper–dataset–paper-baseline interaction chains and uses them to construct reasoning chains. To stimulate reasoning ability, We finetune a language model to generate these chains for reranking, producing interpretable judgments that leverage the interaction-chain prior and go beyond simple semantic matching. This two-stage pipeline couples collective perception–enhanced retrieval with interaction chain–driven reasoning, yielding recommendations that are both semantically grounded and aligned with evaluation intent.
@@
We validate the proposed framework on AgentExpt, our new corpus linking papers (2015–2025, ten flagship AI venues) to baselines and datasets. We evaluate two tasks—baseline and dataset recommendation—under Recall@$k$ and HitRate@$k$. Compared with the strongest prior baseline , our method achieves averaged improvements across the two tasks of +5.85% in Recall@20, +8.30% in HitRate@5, and +7.90% in HitRate@10. Ablations show that collective perception contributes the largest share of retrieval gains, while interaction-chain reasoning drive consistent improvements in hitrate.
The key contributions of this work are as follows:
  • We construct, to our knowledge, the first large-scale, high-quality dataset for baseline and dataset recommendation. Our dataset covers roughly 85% of the baselines and datasets used at top AI conferences over the past five years.
  • We propose a collective perception–enhanced retriever that represents each candidate by jointly encoding its self-description and the community’s perception of how it is positioned and used, yielding retrieval signals that go beyond surface-level textual similarity and better reflect the evaluation intent.
  • We develop a reasoning-augmented reranking framework that leverages interaction chains to construct explicit reasoning chains and finetunes a large language model to produce interpretable justifications and refined rankings.
  • We empirically demonstrate state-of-the-art performance on both baseline and dataset recommendation: averaged over the two tasks, HitRate@5 improves by 8.30%, HitRate@10 by 7.90%, and Recall@20 by 5.85% over the strongest prior baseline.
2. Related Work
2.1. LLM-Based Agents for Research Assistance
Recent efforts have explored the use of LLM-based agents to automate scientific research, ranging from idea generation to full paper drafting. Systems such as ResearchAgent (Baek et al., 2024), IdeaSynth (Pu et al., 2025), and NovelSeek (Team et al., 2025) focus on generating and refining research ideas with literature grounding and structured feedback. In parallel, a line of research targets algorithm optimization itself, where agents search for improved procedures or code, including AlphaEvolve (Novikov et al., 2025), FunSearch (Fawzi and Paredes, 2023). Building on this, end-to-end frameworks including The AI Scientist (Lu et al., 2024; Yamada et al., 2025), Zochi (Intology, 2025), Carl (Institute, 2025), AI-Researcher (Tang et al., 2025), Agent Laboratory (Schmidgall et al., 2025), and domain-specific agents such as Biomni (Huang et al., 2025) attempt to close the entire loop by combining literature review, idea generation, and paper writing. While these systems represent significant progress toward autonomous scientific discovery, they overlook a crucial step in practical research, namely the selection of appropriate baselines and datasets.
2.2. Benchmarks for Baseline and Dataset Recommendation
Early work relied on general dataset search engines that index web metadata and support keyword queries, but they lack task-specific recommendation benchmarks tied to research problem descriptions or scientific contexts (Brickley et al., 2019; Chapman et al., 2020). For baseline recommendation, RecBaselines2023 is, to our knowledge, the only dataset that recommends baselines conditioned on a paper, but its scope is confined to the recommender systems domain and its scale is limited to on the order of several hundred papers and baselines (Ivanova et al., 2023). For dataset recommendation, Färber and Leisinger (2021c) cast the task as multi label classification from natural language problem statements to a fixed inventory of datasets, which simplifies evaluation but cannot naturally handle unseen datasets at test time; Viswanathan et al. (2023a) instead formulate an open domain retrieval benchmark with a large automatically constructed training set and a smaller expert annotated test set, showing that bi encoder retrieval substantially outperforms keyword search and third party engines. However, both lines of work primarily source candidate datasets from portals such as Papers with Code, OpenAIRE, and Wikidata, which means many datasets actually used in published top venue papers but not indexed by these portals are omitted, reducing coverage and limiting practical utility for real research workflows.
2.3. Methods for Baseline and Dataset Recommendation
Prior work spans three main paradigms. A first line treats dataset recommendation as supervised text classification from problem descriptions to a fixed label set of datasets, which is simple and effective on closed inventories but struggles with unseen labels and label drift at test time (Färber and Leisinger, 2021c, b). A second line builds heterogeneous graphs over papers, methods, and datasets and learns links for recommendation, for example via variational graph autoencoders; these approaches leverage co usage and citation topology but face cold start issues and are less natural for free form queries (Qayyum et al., 2025; Altaf et al., 2019). A third line adopts neural retrieval with dense encoders, where a bi encoder retrieves datasets for natural language ideas at open world scale and typically outperforms keyword search, though most systems optimize semantic similarity rather than experimental suitability and offer limited explainability (Viswanathan et al., 2023a). For baseline recommendation, collaborative filtering over paper baseline co occurrences has been shown effective but is domain bound and provides little semantic justification for new ideas (Ivanova et al., 2023).
@@ 2.3. Methods for Baseline and Dataset Recommendation
Figure 3. Pipeline for constructing the AgentExpt knowledge base. We (1) download and parse papers (from flagship AI conferences), (2) identify baselines and datasets by locating experiment sections, (3) apply rule-based and LLM-based filtering using citation frequency, naming consistency, and contextual positioning to prune false positives. The final dataset contains 108,825 papers, 116,970 baseline entities, 68,316 dataset entities, and their respective cross‑entity connections.:
3. Preliminary
Table 1. Comparison of datasets for recommending baselines or datasets.
@@ 3.2. Input Representation
$q\;=\;\mathrm{Summ}\!\big(a(p)\big),$
where $\mathrm{Summ}(\cdot)$ is implemented with GPT4o (OpenAI, 2024) to produce a concise synopsis that preserves the problem setting, assumptions, and evaluation intent. This abstract based representation captures the key scientific context.
3.3. Evaluation Metrics
@@
  • HitRate@$k$: A binary measure that equals 1 if at least one relevant item is retrieved within the top-$k$, and 0 otherwise. This reflects a user-centric view of success: whether the system surfaces some useful component early.
4. Data Collection
4.1. Overview
@@
To effectively support the retrieval and recommendation of state-of-the-art baselines and suitable evaluation datasets corresponding to specific research topics, it is imperative to precisely extract baselines and datasets utilized in empirical evaluations reported in scholarly literature. Accomplishing this task requires extensive extraction of structured entities accompanied by comprehensive metadata associated with respective papers, baselines, and datasets. However, current academic resources (e.g., Papers with Code(Pap, [n. d.]), Hugging Face(Hug, [n. d.])) and automated information extraction methodologies (e.g., DataFinder(Viswanathan et al., 2023b), DataHunter(Färber and Leisinger, 2021a)) exhibit substantial limitations concerning dataset coverage comprehensiveness, entity metadata completeness, and the accuracy of identifying PaperBaseline, PaperDataset associations.
@@
Specifically, establishing a robust and high-quality Paper-Baseline-Dataset data involves addressing two primary challenges. First, significant heterogeneity in data sources and formats complicates information extraction efforts. Baselines often appear in abbreviated or inconsistent terminologies across different publications, while datasets frequently suffer from ambiguous naming conventions and inconsistent references. Moreover, the absence of explicit links within papers to original baseline implementations and dataset repositories further complicates automated text parsing, entity recognition, and information alignment processes. Second, the exponential growth of artificial intelligence research outputs exacerbates inefficiencies and scalability constraints inherent to manual or semi-automated annotation methods. Leading AI conferences receive tens of thousands of submissions annually, making traditional annotation approaches impractical and highlighting the need for scalable automation.
@@
Motivated by recent advancements in automated scientific information extraction methodologies such as DataFinder (Viswanathan et al., 2023b), we have developed and implemented a comprehensive, end-to-end automated pipeline for scholarly data extraction and validation. Our pipeline specifically targets research papers published between 2015 and 2025 across ten prominent artificial intelligence conferences: NeurIPS (Neu, [n. d.]), ICML (ICM, [n. d.]), ICLR (ICL, [n. d.]), ACL (ACL, [n. d.]), NAACL (NAA, [n. d.]), EMNLP (EMN, [n. d.]), CVPR (CVP, [n. d.]), ICCV (ICC, [n. d.]), AAAI (AAA, [n. d.]), and IJCAI (IJC, [n. d.]). As a result, we have constructed a scholarly knowledge repository distinguished by extensive data coverage and remarkable accuracy.
4.2. Automated Data Collection Pipeline
The proposed automated pipeline comprises critical stages including paper acquisition and metadata extraction, structured content parsing, candidate entity recognition for baseline methods and datasets, and a rigorous validation process employing both rule-based filtering and large language model (LLM)-based verification. This pipeline significantly enhances entity extraction precision and recall. Furthermore, it systematically integrates essential metadata attributes for each entity, such as repository links, publication year, original data sources, evidence supporting entity recognition, and standardized canonical naming conventions. The end-to-end automation effectively mitigates challenges associated with data heterogeneity and substantially alleviates scalability issues encountered by manual annotation methodologies, thereby providing robust, accurate, and valuable data for academic recommendation and literature retrieval tasks.
Figure 4. The coverage of experiment baselines and datasets. The vertical axis indicates the coverage, calculated as the fraction of resources employed in year N that were introduced in preceding years (1 to N-1), reflecting the dependency on established experimental components over time.:

### v3

@@
Abstract
In modern AI research, baseline and dataset selection is a high-stakes decision in experimental design. It operationalizes a research idea into a concrete evaluation protocol and largely determines the validity and comparability of empirical conclusions. However, making appropriate choices is increasingly difficult as baselines and datasets proliferate, while suitability is inherently context-dependent and rarely captured by baseline and dataset metadata. To address these challenges, we present AgentExpt, a comprehensive framework for baseline and dataset recommendation. We first curate a large-scale, high-quality knowledge base that links 108,825 accepted papers to their used baselines and datasets. Based on this resource, we design a collective perception-enhanced retriever that represents each baseline or dataset by integrating first-person self-descriptions with third-person citation contexts, thereby effectively positioning them within the scholarly network. We further design a reasoning-augmented reranker that encodes baseline-dataset interaction chains as a reasoning prior to fine-tune an LLM, producing refined rankings with interpretable justifications. Experiments show that our framework outperforms the strongest baseline, with average gains of +5.85% in Recall@20 and +7.90% in HitRate@10, and ablation studies confirm the effectiveness of our designed components. Overall, AgentExpt advances the efficient and reliable automation of experimental design.
Machine Learning, ICML
Figure 1: Overview of the research problem and our proposed solution.
1 Introduction
Baseline and dataset selection is a recurring bottleneck in modern AI research. Beyond being a clerical choice, it determines the evaluation target, the comparison standard, and the data distribution under which claims are tested, thereby shaping whether results are comparable, reproducible, and scientifically meaningful (Torralba and Efros, 2011). This step has become increasingly difficult as the ecosystem of baselines and datasets expands and fragments across tasks, versions, and evaluation conventions (Ott et al., 2022). Crucially, experimental suitability is context-dependent: a baseline or dataset that is appropriate in one setting may be misleading or uninformative in another (Paullada et al., 2021), yet such suitability cues are rarely explicit in item metadata or portal descriptions (Figure 1).
In parallel, recent progress in large language models (LLMs) and agentic systems has spurred interest in automated research workflows (Brown et al., 2020; Radford et al., 2019, 2018; Shang et al., 2024b, a; Li et al., 2025; Zhao et al., 2025), ranging from literature review and idea generation (Huang et al., 2025; Team et al., 2025) to broader pipeline automation (Institute, 2025; Intology, 2025; Lu et al., 2024; Yamada et al., 2025). However, most existing systems focus on upstream stages or operate in controlled experimental settings (e.g., fixed benchmarks and predefined tasks), and therefore treat baseline and dataset choices as given rather than as a decision variable. As a result, reliable automation of baseline and dataset recommendation remains comparatively underexplored despite being central to rigorous experimental design.
Three limitations in existing research remain salient. First, coverage is limited: many systems draw candidates from domain-restricted collections or portal-curated inventories (e.g., Paper With Code (n.d.)), which do not reliably reflect what is actually used across the literature and therefore miss a substantial portion of relevant baselines and datasets (Ivanova et al., 2023; Viswanathan et al., 2023; Färber and Leisinger, 2021a). Second, suitability is under-modeled: even when using supervised text classification (Färber and Leisinger, 2021b, a), dense retrieval (Viswanathan et al., 2023), graph-based models (Altaf et al., 2019; Qayyum et al., 2025), or collaborative filtering (Ivanova et al., 2023), most approaches rely on first-person descriptions and metadata, effectively treating suitability as topical similarity and overlooking how artifacts are positioned and used in practice. Third, baseline selection and dataset selection are typically decoupled, and thus recommendations ignore their joint compatibility and the fact that evaluation intent is often realized through baseline-dataset pairings rather than either component in isolation.
@@ 1 Introduction
In contrast, day-to-day research practice typically starts from a small set of closely related papers and follows their trails to the baselines and datasets they actually used, rather than browsing portal listings; researchers also consider baselines and datasets jointly, favoring pairings that have been used together or are methodologically compatible with the intended setting. Moreover, experimental choices must be justified by situating selected baselines and datasets relative to prior work, making citation contexts a valuable third-person signal that complements first-person self-descriptions. Together, these observations motivate a unified formulation that links papers, baselines, and datasets and explicitly leverages community-level perception.
Motivated by these observations, we present AgentExpt, a comprehensive framework for baseline and dataset recommendation that directly targets the three limitations above (Figure 1). To address limited coverage, we curate a large-scale, literature-grounded knowledge base that links papers to the baselines and datasets they actually use. Concretely, we build an automated data-collection pipeline that gathers 108,825 accepted papers from the past decade across ten flagship AI venues, parses the experiment sections from PDFs, and employs LLM-assisted extraction to identify baseline and dataset mentions. We then perform entity normalization and source linking to resolve naming variations and map extracted mentions to canonical components, yielding a high-precision paper–component linkage that reflects real usage in experiments rather than portal listings.
To model context-dependent suitability beyond metadata-only similarity, we design a collective perception-enhanced retriever that augments first-person self-descriptions with third-person evidence from the literature (Figure 4). Specifically, for each baseline or dataset, we extract its citation contexts from downstream papers as community-level perception signals, summarize these contexts into an aggregate usage profile, and integrate it with artifact’s self-description to form a dual-view representation. We then fine-tune an embedding model on these representations to efficiently retrieve candidates whose usage patterns, not only topical wording, align with the evaluation intent implied by a query.
Finally, to capture the synergy between baselines and datasets, we design a reasoning-augmented reranker that leverages paper-mediated interaction evidence between baselines and datasets to construct structured reasoning traces. This module extracts paper-baseline-paper-dataset (P-B-P-D) and paper-dataset-paper-baseline (P-D-P-B) interaction chains, and uses them to construct reasoning chains. We use these traces as a reasoning prior to fine-tune an LLM for reranking, enabling refined recommendations accompanied by interpretable justifications that are both semantically grounded and aligned with evaluation intent.
@@
We evaluate AgentExpt on the curated knowledge base against four families of baselines. Compared with the strongest baseline, AgentExpt yields average gains (across the two tasks) of +5.85% in Recall@20, +8.30% in HitRate@5, and +7.90% in HitRate@10. Ablations show that collective perception contributes the largest share of retrieval gains, while interaction-chain reasoning drives consistent improvements in hit rates.
We summarize the key contributions as follows:
  • We present AgentExpt, a comprehensive framework that automates baseline and dataset recommendation for experimental design under context-dependent suitability.
  • We curate a large-scale, high-quality knowledge base that links 108,825 papers to the baselines and datasets they use, enabling realistic training and evaluation for AgentExpt and future research.
  • We design a collective perception-enhanced retriever, integrating first-person self-descriptions with third-person citation contexts to capture community usage signals beyond metadata-only matching.
  • We design a reasoning-augmented reranker, encoding baseline-dataset interaction chains as a reasoning prior to fine-tune an LLM for refined rankings with interpretable justifications.
  • We empirically validate the state-of-the-art performance of AgentExpt, and the effectiveness of each component.
2 Related Work
2.1 LLM-Based Agents for Research Assistance
LLM-based agents have increasingly automated scientific workflows, ranging from idea generation (Baek et al., 2024; Pu et al., 2025; Team et al., 2025) and algorithm optimization (Novikov et al., 2025; Fawzi and Paredes, 2023) to end-to-end paper production (Lu et al., 2024; Yamada et al., 2025; Intology, 2025; Institute, 2025; Tang et al., 2025; Schmidgall et al., 2025; Huang et al., 2025). However, these systems generally bypass detailed experimental design, failing to systematically recommend appropriate baselines and datasets essential for rigorous evaluation. In contrast, we introduce a framework to automate baseline and dataset retrieval.
2.2 Datasets for Baseline and Dataset Recommendation
Early search engines lacked scientific context awareness (Brickley et al., 2019; Chapman et al., 2020), while the only paper-conditioned baseline recommendation dataset, RecBaselines2023 (Ivanova et al., 2023), is limited in scale and restricted to recommender systems. For dataset recommendation, approaches have evolved from fixed-label classification (Färber and Leisinger, 2021b) to open-domain retrieval (Viswanathan et al., 2023). However, these datasets typically source candidates from public portals (e.g., Wikidata, Papers with Code) rather than actual literature, resulting in significant coverage gaps regarding artifacts used in published experiments. In contrast, we construct a comprehensive dataset by extracting baselines and datasets directly from papers across ten flagship AI venues, ensuring high coverage of real-world research practices.
2.3 Methods for Baseline and Dataset Recommendation
Prior approaches span supervised classification to fixed labels (Färber and Leisinger, 2021b, a), graph-based learning on heterogeneous networks (Qayyum et al., 2025; Altaf et al., 2019), and dense neural retrieval (Viswanathan et al., 2023), alongside collaborative filtering strategies for baselines (Ivanova et al., 2023). However, these methods typically prioritize surface-level semantic similarity or structural co-occurrence, often overlooking experimental suitability and failing to provide interpretable justifications for their choices. In contrast, we propose a framework that integrates collective citation perception to capture usage intent and leverages interaction-chain reasoning to deliver recommendations that are both experimentally suitable and interpretable.
@@ 2.3 Methods for Baseline and Dataset Recommendation
Figure 2: Pipeline for constructing the AgentExpt knowledge base. We (1) download and parse papers (from flagship AI conferences), (2) identify baselines and datasets by locating experiment sections, (3) apply rule-based and LLM-based filtering using citation frequency, naming consistency, and contextual positioning to prune false positives. The final dataset contains 108,825 papers, 116,970 baseline entities, 68,316 dataset entities, and their respective cross‑entity connections.
3 Preliminary
Table 1: Comparison of datasets for recommending baselines or datasets.
@@ 3.2 Input Representation
$q\;=\;\mathrm{Summ}~\!\big(a(p)\big),$
where $\mathrm{Summ}(\cdot)$ is implemented with GPT-4o (OpenAI, 2024) to produce a concise synopsis that preserves the problem setting, assumptions, and evaluation intent. This abstract based representation captures the key scientific context.
3.3 Evaluation Metrics
@@
  • HitRate@$k$: A binary measure that equals 1 if at least one relevant item is retrieved within the top-$k$, and 0 otherwise. This reflects a user-centric view of success: whether system surfaces some useful component early.
4 Dataset Construction
4.1 Overview
@@
To effectively support the retrieval and recommendation of state-of-the-art baselines and suitable evaluation datasets corresponding to specific research topics, it is imperative to precisely extract baselines and datasets utilized in empirical evaluations reported in scholarly literature. Accomplishing this task requires extensive extraction of structured entities accompanied by comprehensive metadata associated with respective papers, baselines, and datasets. However, current academic resources (e.g., Paper With Code (n.d.), HuggingFace (n.d.)) and automated information extraction methodologies (Viswanathan et al., 2023; Färber and Leisinger, 2021a)) exhibit substantial limitations concerning dataset coverage, entity metadata completeness, and the accuracy of identifying Paper-Baseline, Paper-Dataset associations.
@@
Specifically, establishing a robust and high-quality Paper-Baseline-Dataset data involves addressing two primary challenges. First, significant heterogeneity in data sources and formats complicates information extraction efforts. Baselines often appear in abbreviated or inconsistent terminologies across different publications, while datasets frequently suffer from ambiguous naming conventions and inconsistent references. Moreover, the absence of explicit links within papers to original baseline implementations and dataset repositories further complicates automated text parsing, entity recognition, and information alignment processes. Second, the exponential growth of research outputs exacerbates inefficiencies and scalability constraints inherent to manual or semi-automated annotation methods. Leading AI conferences receive tens of thousands of submissions annually, making traditional annotation approaches impractical and highlighting the need for scalable automation.
@@
Motivated by recent advancements in automated scientific information extraction (Viswanathan et al., 2023), we design a comprehensive, end-to-end automated pipeline for scholarly data extraction and validation. Our pipeline specifically targets research papers published between 2015 and 2025 across ten prominent AI conferences: NeurIPS (n.d.), ICML (n.d.), ICLR (n.d.), ACL (n.d.), NAACL (n.d.), EMNLP (n.d.), CVPR (n.d.), ICCV (n.d.), AAAI (n.d.), and IJCAI (n.d.), because they collectively represent flagship venues spanning major AI subfields with consistently high review standards, providing a literature-grounded and high-quality source for extracting experimental configurations at scale.
4.2 Automated Dataset Construction Pipeline
To address the challenges of data heterogeneity and the scalability limitations of manual annotation, we design an end-to-end automated pipeline. Specifically, the pipeline comprises three critical stages: Acquisition and Parsing, Baseline and Dataset Identification, and Rule-based and LLM-based Filtering, see Figure  2.